if the reservoir pipe feeds the juice into the tank of diameter 2m calculate how long it will take to fill this tank with the juice to a height of 50 cm


1
Expert's answer
2021-08-16T17:30:47-0400

Gives

d=2m

h=50cm=0.5m

Water speed=v

t=?

"t=\frac{\pi r^2h}{v}"

Put value

"t=\frac{3.14 \times 1\times0.5}{v}"


"t=\frac{1.57}{v}sec"